OpenStudy (anonymous):

Given the functions f (x) = x +1 and g(x) = 3x − 2x^2 , find: f (f (-3))

OpenStudy (anonymous):

-2

OpenStudy (agent0smith):

f (f (-3)) this means you first find f(-3)... which means you need to plug -3 in place of x into: x +1 Whatever that is equal to, you will plug that number into x+1 again. The result of that is f(f(-3))

OpenStudy (anonymous):

I got -2

OpenStudy (agent0smith):

Show your work.

OpenStudy (agent0smith):

f(-3) = -3 +1 then you still have to plug that again into x+1

OpenStudy (anonymous):

f (x) = x +1 = f(x)=(-3+1)= -2 -2+1 = -1

OpenStudy (agent0smith):

Good job.

OpenStudy (anonymous):

okay thanks so I have to solve and then plug back in

OpenStudy (agent0smith):

Yes, f(-3) is -2. f(f(-3)) is thus the same as f(-2) and f(-2) = -2 + 1.

OpenStudy (anonymous):

if I have: Given the functions f(x)=2x/x+5 and g(x)= 7-x/x-1 and I have to find f (g (3)) I plug into - 7-3/3-1 = 4/2 = 2 and then plug 2 back into 7-2/2-1 = 5/1 ?

OpenStudy (anonymous):

no ... that's not right

OpenStudy (agent0smith):

You're on the right track. For f( g(3) ) - This time find g(3), which you've done. Now plug that number into f(x).

OpenStudy (anonymous):

f(x) = -(4/3)

OpenStudy (agent0smith):

if f(x)=2x/x+5 where did you get -4/3 from? f( g(3) ) means f(2), since g(3) = 2 as you found.

OpenStudy (anonymous):

oh! g(x)=4/7

OpenStudy (agent0smith):

Excellent.

OpenStudy (anonymous):

so if it asks me for f(g) I plug that into f(x) for x and then that answer I plug it into g(x) for x and vice versa. =)

OpenStudy (agent0smith):

Yep, basically you just start from inside the brackets, and move towards the outside eg for a more complex example, f( g( f(3) ) ) -start from the very inside part for that, you'd first find f(3). Then plug that number into g(x). Then plug that number into f(x).

OpenStudy (anonymous):

(a) f (g(1)) (b) g( f (1)) (c) f (g(x)) (d) g( f (x)) (e) f ( f (1)) (f) g(g(1)) (g) f ( f (x)) (h) g(g(x)) what does the f(f(x)) mean and the same for the g(g(x)) ? I understand f(g) but not f(f)

OpenStudy (agent0smith):

Always just start from the inside, doesn't matter the letters, same process. f( f(3) ) means find f(3) then plug that result into f. g (g-2) ) means find g(-2) and plug that back into g.

OpenStudy (anonymous):

ok thanks =)

OpenStudy (anonymous):

one more questions?

OpenStudy (agent0smith):

sure

OpenStudy (agent0smith):

Oh sorry i misread you above. f( f(x) ) means you have to plug the whole thing in, in place of x. If f(x) = 2x + 3 then f( f(x) ) is f(2x+3) = 2(2x+3) + 3 ^see, just replace the x, with the actual function 2x+3

OpenStudy (anonymous):

lets say this is my problem:Given the functions f (x) = x^2 + 2 and g(x) = 5x −8 , find: (a) f (g(1)) (c) f (g(x)) (e) f ( f (1)) (g) f ( f (x)) I understand the f(g(1)) - plug that into f(x) and the answer I get from that plug it into g(x). Would I plug the f(f(1)) into f(x) again even though I did that for A already?A

OpenStudy (agent0smith):

f( g(1) ) means find g(1) and plug that into f(x) f( f(1) ) means find f(1) and plug that into f(x)

OpenStudy (agent0smith):

f ( g(x) ) means you need to plug the whole g(x), which is (5x −8) in place of x into f(x)

OpenStudy (anonymous):

would I take that answer and put it back into g(x)?

OpenStudy (agent0smith):

Which?

OpenStudy (anonymous):

f(g(x) plug (5x-8) in place of x into f(x)

OpenStudy (agent0smith):

Yes.